A soccer field has a rectangular perimeter of 230 feet. The length of the field is twice as long as the width. What is the length of each side and width of each side?
Answer by TimothyLamb(4379)

2L + 2w = 230
L = 2w
---
2L + 2w = 230
2(2w) + 2w = 230
4w + 2w = 230
6w = 230
---
w = 38.33333 ft
L = 76.66666 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php